OpenStudy (anonymous):

x^2 – 36 = 0 and x^2 = 8x – 12 Part 2: Describe what the solution(s) represent to the graph of each.

OpenStudy (anonymous):

x^2 - 36 = 0. x = 6 or -6

OpenStudy (anonymous):

x^2 = 8x – 12 x =6 and x = 2.

OpenStudy (anonymous):

Now... Describe what the solution(s) represent to the graph of each.

OpenStudy (anonymous):

so basically. when you find the solutions to a function, you've found the x-intercepts, or in other words, when the function crosses the x-axis.

OpenStudy (anonymous):

That's what my answer would be?

OpenStudy (anonymous):

I'm guessing so. I mean that's the general answer for a question like that. I'm not sure whether that's precisely what you're looking for though.

OpenStudy (anonymous):

It'll do. I have one more question on this Problem.